# mode = title

### v1

@@ 3.2 Monocular Depth Estimation Module (MDEM)
To address the challenges in monocular depth estimation, particularly generalization across diverse datasets, our MDEM employs the innovative architecture of Zoe [35]. Zoe model utilizes a distinctive two-stage framework:
@@
    Pre-training on a wide array of datasets for relative depth estimation, which fosters excellent generalization ability.
@@ 4 Materials
4.1 Training Datasets for CyclePose
The training and validation processes were performed on the CyclePose module because it was the only model we trained in this study. Two distinct datasets were utilized - an extensive internal dataset for training and selected subsets from the EndoSLAM dataset [16] for validation. For training, we leveraged a large internal unlabeled dataset comprising more than 300 hours of gastroscopy and prostatectomy videos, collected from 100 patients undergoing these procedures at Fondazione Policlinico Universitario Campus Bio-Medico. In total, this training dataset contained 2,250,900 frames, allowing us to effectively train our pose estimation network on a substantial amount of endoscopic data. To prepare the training data, we processed the videos by extracting individual frames and performing a center crop of size 128x128 pixels on each frame. This center cropping approach enhances system robustness by training it to work effectively with a reduced receptive field, while avoiding any distortion that may arise from resizing the frames. By focusing on the central region of each frame, we ensure that the most relevant visual information is preserved while maintaining the spatial integrity of the data. To validate CyclePose, we used the following subsets from the EndoSLAM dataset, which are completely separate from the training data: HighCam Colon IV Tumor-Free Trajectory 1, LowCam Colon IV Tumor-Free Trajectory 1, HighCam Colon IV Tumor-Free Trajectory 5, and LowCam Colon IV Tumor-Free Trajectory 5. These subsets account for four video sequences with a total of 3024 frames, providing a diverse range of scenarios to assess CyclePose’s inference performance. Importantly, there is no overlap between the training and validation datasets, as the EndoSLAM subsets were specifically chosen for validation purposes only and were not used in any way during the training phase.
@@ 4.2 Testing Datasets for Depth and Pose Estimation Modules
To evaluate the performance of our depth estimation and pose estimation modules we utilized three datasets with distinct characteristics, as summarized in Table 1. They are:
@@
    EndoSLAM Dataset [16]: In addition to its use in validating CyclePose, the EndoSLAM dataset was also employed for testing the pose estimation module of our method. The subsets used for validating CyclePose were excluded from the testing process to maintain a clear separation between validation and testing data. While the EndoSLAM dataset includes depth data, we did not use it for testing depth estimation due to its synthetic nature. The dataset comprises both ex-vivo and synthetic porcine data, with 35 sub-datasets totaling 76,837 frames spanning various gastrointestinal areas, providing pose data for a comprehensive evaluation of our method’s performance in pose estimation.
@@ 5 Experimenental Configuration
To validate the effectiveness and generalization capabilities of our proposed MVSLAM approach, we conducted a comprehensive comparative analysis against two state-of-the-art approaches [1] in the field of endoscopic surgery: EndoSLAM [16] and EndoDepth [17]. These methods were selected based on their proven performance and distinct approaches to addressing the challenges inherent in endoscopic imagery. EndoSLAM, also known as EndoSfMLearner, is an unsupervised monocular visual odometry and depth estimation approach designed specifically for endoscopic videos. It focuses on depth estimation, pose estimation, and 3D reconstruction. The depth estimation component, based on [40], predicts dense disparity maps from single images using spatial attention mechanisms and a hybrid loss function. The pose estimation component calculates the relative 6-DoF camera poses between consecutive image frames, leveraging a spatial attention block (ESAB) to emphasize texture details and depth differences.
EndoDepth presents a comprehensive approach for 3D reconstruction and camera motion estimation in monocular endoscopic sequences. It utilizes self-supervised depth networks, photometric tracking, and volumetric fusion to generate pseudo-RGBD frames, track the camera pose, and fuse registered depth maps into a coherent 3D scene model. The depth estimation component employs the Monodepth2 [33] network architecture, which is self-supervised and operates on a U-Net encoder-decoder structure with a ResNet18 encoder. Both EndoSLAM and EndoDepth were utilized as developed by their respective authors, without any modifications or retraining. The training and validation contexts of these models are also significant. EndoDepth validation on the Hamlyn Dataset and EndoSLAM on the EndoSLAM dataset provide a varied baseline for assessing the generalization capabilities of our proposed MVSLAM approach.
@@ 5.1 Performance Metrics and Statistical Analysis
To evaluate the performance of the SLAM approach, we employed a set of well-established metrics for depth estimation and pose estimation. For depth estimation, we selected the Absolute Relative Difference (Abs, Rel. Diff), Root Mean Square Error (RMSE), RMSE log, Squared Relative Error (Sq. Rel.), and accuracy thresholds of $1.25$, $1.25^{2}$, and $1.25^{3}$, as described in [41] and presented in Table 2.
@@
To measure the pose estimation performance, we utilized the Relative Translation Error (RTE), Relative Rotation Error (RRE), and Absolute Translation Error (ATE), as shown in Table 3.
@@
Table 3: Trajectory Evaluation Metrics (evo Implementation). $E_{ij}$: pose error between estimated and ground truth poses. $\text{trans}(E_{ij})$: translational component, $\text{angle}(\cdot)$: angular difference from rotational data. ATE: Absolute Trajectory Error, RTE: Relative Trajectory Error, RRE: Relative Rotation Error.
These metrics are widely used in the literature for evaluating the accuracy of pose estimation in SLAM systems [1]. In addition to the performance metrics, we conducted a statistical analysis to determine the significance of the performance differences observed between the MVSLAM approach and the benchmark models across various scenarios. We verified that the data for each model meets the assumptions required for using an independent two-sample T-test, including normality and equal variances. We then performed pairwise comparisons between the MVSLAM approach and each benchmark model separately, using a significance level of $p=0.05$.
@@ 6 Results
In this section, we present the results of our comprehensive evaluation aimed at validating the efficacy and robustness of our proposed MVSLAM approach. Initially, in section 6.1, we provide an analysis of the monocular pose estimation module. This is followed by section 6.2, where we explore the application and validation of the depth estimation module. It should be noted that the 3D Mapping Module (3DM) cannot be directly evaluated due to two key factors. First, the lack of ground truth 3D models for the tested environments hinders a comprehensive validation of the reconstruction quality. Second, even if ground truth 3D models were available, the scale ambiguity inherent in the monocular pose and depth estimation modules would render the reconstructed model incomparable due to differing scales. However, this inability to directly evaluate the 3DM does not represent a limitation of our approach. The 3DM employs state-of-the-art 3D reconstruction algorithms that have been rigorously tested and validated in prior work [39, 42]. Furthermore, the 3D mapping and localization capabilities critically depend on the performance of the pose estimation and depth estimation modules, which are thoroughly evaluated in the following sections. The successful functioning of the overall SLAM system thus implicitly validates the efficacy of the 3DM.
@@ 6.1 Monocular Pose Estimation Module (MPEM)
Figure 4 presents the performances of the monocular pose estimation module measured in terms of ATE, RTE, and RRE on the two datasets (SCARED and EndoSLAM) when adopting each of the three methods, i.e., CyclePose, EndoSfmLearner and EndoDepth.
Figure 4: Comparison of the performance of CyclePose, EndoDepth, and EndoSfmLearner algorithms on the SCARED and ENDOSLAM datasets. Metrics evaluated include Absolute Trajectory Error (ATE), Relative Trajectory Error (RTE), and Relative Rotation Error (RRE). Each box plot shows the distribution of the errors for the respective algorithms, highlighting the median, interquartile range, and outliers.
In the case of trajectory errors, being either an absolute or a relative error, both rows show that CyclePose and EndoSfmLearner have lower median scores and interquartile rages than EndoDepth. The figure compares the three methods across the two datasets and three metrics. The rows represent the SCARED and ENDOSLAM datasets, while the columns show the ATE, RTE, and RRE metrics respectively. Regarding the statistical analysis, Table 4 shows there none of these methods has a performance that is statistically different from the other two. It is also interesting to note that our CyclePose approach not only performs as the others but it also has the lowest inference time (Table 5).
DS Metric Comparison T-Stat P-Value
@@
BodySLAM vs EndoSFMLearner 0.63 0.5334
@@
ENDOSLAM ATE BodySLAM vs EndoDepth -1.02 0.3121
BodySLAM vs EndoSFMLearner -0.12 0.9087
@@
BodySLAM vs EndoSFMLearner -0.01 0.9941
@@
Table 4: Comparative analysis of BodySLAM against EndoDepth and EndoSFMLearner across different metrics (ATE, RTE, RRE) for SCARED and ENDOSLAM datasets. T-Statistic and P-Value are reported for each comparison. Bold P-Values indicate statistically significant differences
(p < 0.05).
Model Dataset Time [s]
@@
CyclePose EndoSLAM 3.32
@@
Absolute Relative Difference BodySLAM vs EndoDepth -2.18 0.0362
BodySLAM vs EndoSFMLearner -4.24 0.0001
RMSE BodySLAM vs EndoDepth -1.99 0.0536
@@
RMSE Log BodySLAM vs EndoDepth -2.13 0.0399
BodySLAM vs EndoSFMLearner -4.42 8.71e-05
Squared Relative Error BodySLAM vs EndoDepth -2.15 0.0381
BodySLAM vs EndoSFMLearner -9.69 1.44e-11
Accuracy (1.25) BodySLAM vs EndoDepth 5.14 9.69e-06
BodySLAM vs EndoSFMLearner 2.55 0.0150
Accuracy (1.25)² BodySLAM vs EndoDepth 4.12 0.0002
BodySLAM vs EndoSFMLearner 1.54 0.1314
Accuracy (1.25)³ BodySLAM vs EndoDepth 3.56 0.0011
BodySLAM vs EndoSFMLearner 1.37 0.1803
Table 6: Statistical Performance Analysis of the Monocular Depth Estimation Module on the Hamlyn Dataset. The table details outcomes for Absolute Relative Difference, Root Mean Square Error (RMSE), RMSE Log, Squared Relative Error, and accuracy thresholds ($1.25$, $1.25^{2}$, $1.25^{3}$). Results include pairwise comparisons of BodySLAM versus EndoDepth and BodySLAM versus EndoSFMLearner, assessed through t-tests with a significance level of
p < 0.05.
@@
Absolute Relative Difference BodySLAM vs EndoDepth -9.69 1.44e-11
BodySLAM vs EndoSFMLearner -4.14 0.0002
RMSE BodySLAM vs EndoDepth -1.01 0.3211
@@
RMSE Log BodySLAM vs EndoDepth -14.99 4.73e-17
BodySLAM vs EndoSFMLearner -8.35 6.04e-10
Squared Relative Error BodySLAM vs EndoDepth -0.85 0.4018
BodySLAM vs EndoSFMLearner 0.31 0.7598
Accuracy (1.25) BodySLAM vs EndoDepth 13.10 2.92e-15
BodySLAM vs EndoSFMLearner 5.82 1.22e-06
Accuracy (1.25)² BodySLAM vs EndoDepth 10.13 4.37e-12
BodySLAM vs EndoSFMLearner 2.56 0.0147
Accuracy (1.25)³ BodySLAM vs EndoDepth 8.53 3.65e-10
BodySLAM vs EndoSFMLearner 1.26 0.2151
Table 7: Statistical Performance Analysis of the Monocular Depth Estimation Module on the SCARED Dataset. The table details outcomes for Absolute Relative Difference, Root Mean Square Error (RMSE), RMSE Log, Squared Relative Error, and accuracy thresholds ($1.25$, $1.25^{2}$, $1.25^{3}$). Results include pairwise comparisons of BodySLAM versus EndoDepth and BodySLAM versus EndoSFMLearner, assessed through t-tests with a significance level of
p < 0.05.
Turning our attention to RRE, i.e., a score measuring the angular deviation in the rotation component, we notice that CyclePose outperforms both EndoSfmLearner and EndoDepth on both datasets (both rows and last column of Figure 4). Furthermore, these performance differences are also statistically significant (last rows for both datasets of Figure 4). To deepen the results we also perform an ablation study that aims to investigate the effectiveness of the new terms that we introduced for the cycle loss (section 3.1). To this end, we ran four experiments, two per dataset, setting $\lambda_{2}\neq 0$, $\lambda_{2}=0$ (Figure 5).

### v2

@@ 3.2 Monocular Depth Estimation Module (MDEM): Zoe
To address the challenges of monocular depth estimation in endoscopic environments, particularly the issue of data scarcity, we employ the innovative Zoe architecture [26]. This represents the first use of a well-generalized model for depth prediction in endoscopic SLAM, tackling the data scarcity problem inherent in medical imaging applications.
Zoe model utilizes a distinctive two-stage framework:
@@
    Pre-training on a wide array of datasets for relative depth estimation, which fosters excellent generalization ability. The datasets used for pre-training include.
@@ 4 Materials
4.1 Training Datasets for CycleVO
The training and validation processes were performed on the CycleVO module because it was the only model we trained in this study. Two distinct datasets were utilized: an extensive internal dataset for training and selected subsets from the EndoSLAM dataset [23] for validation. For training, we leveraged a large internal unlabeled dataset comprising more than 300 hours of gastroscopy and prostatectomy videos, collected from 100 patients undergoing these procedures at Fondazione Policlinico Universitario Campus Bio-Medico. In total, this training dataset contained 2,250,900 frames, allowing us to effectively train our pose estimation network on a substantial amount of endoscopic data. To prepare the training data, we processed the videos by extracting individual frames and performing a center crop of size 128x128 pixels on each frame. This center cropping approach enhances system robustness by training it to work effectively with a reduced receptive field, while avoiding any distortion that may arise from resizing the frames. By focusing on the central region of each frame, we ensure that the most relevant visual information is preserved while maintaining the spatial integrity of the data. To validate CycleVO, we used the following subsets from the EndoSLAM dataset, which are completely separate from the training data: HighCam Colon IV Tumor-Free Trajectory 1, LowCam Colon IV Tumor-Free Trajectory 1, HighCam Colon IV Tumor-Free Trajectory 5, and LowCam Colon IV Tumor-Free Trajectory 5. These subsets account for four video sequences with a total of 3024 frames, providing a diverse range of scenarios to assess CycleVO’s inference performance. Importantly, there is no overlap between the training and validation datasets, as the EndoSLAM subsets were specifically chosen for validation purposes only and were not used in any way during the training phase.
@@ 4.2 Testing Datasets for Depth and Pose Estimation Modules
To evaluate the performance of our depth estimation and pose estimation modules we utilized three comprehensive datasets that represent the largest publicly available endoscopic MVSLAM datasets to date. These datasets, summarized in Table 1, provide diverse scenarios and data type to assess the robustness and generalizability of our method:
@@
    EndoSLAM Dataset [23]: In addition to its use in validating CycleVO, the EndoSLAM dataset was also employed for testing the pose estimation module of our method. The subsets used for validating CycleVO were excluded from the testing process to maintain a clear separation between validation and testing data. While the EndoSLAM dataset includes depth data, we did not use it for testing depth estimation due to its synthetic nature. The dataset comprises both ex-vivo and synthetic porcine data, with 35 sub-datasets totaling 76,837 frames spanning various gastrointestinal areas, providing pose data for a comprehensive evaluation of our method’s performance in pose estimation.
@@ 5 Experimenental Configuration
To validate the effectiveness and generalization capabilities of our proposed MVSLAM approach, we conducted a comprehensive comparative analysis against four state-of-the-art approaches in the field of endoscopic MVSLAM and 3D reconstruction.
EndoSLAM [^1] https://github.com/CapsuleEndoscope/EndoSLAM [23], also known as Endo-SfMLearner, is an unsupervised monocular visual odometry and depth estimation approach designed specifically for endoscopic videos. It employs a DepthNet based on a U-Net [40] architecture with ResNet-18 encoder to predict dense disparity maps from single images. The PoseNet estimates six Degree of freedom relative camera poses between consecutive frames. Key innovations include a spatial attention module to focus on highly textured regions and an affine brightness transformer to handle illumination changes.
EndoDepth [^2] https://github.com/UZ-SLAMLab/Endo-Depth-and-Motion [22] presents an approach for 3D reconstruction and camera motion estimation in monocular endoscopic sequences. It utilizes self-supervised depth networks, photometric tracking, and volumetric fusion to generate pseudo-RGBD frames, track camera pose, and fuse depth maps into a 3D model. The depth estimation uses the Monodepth2 [34] network architecture with a ResNet-18 [41] encoder. Camera poses are estimated by minimizing photometric error between frames.
Shao et al. [^3] https://github.com/ShuweiShao/AF-SfMLearner [24] propose a self-supervised framework for monocular depth and ego-motion estimation in endoscopy that introduces the concept of appearance flow to handle brightness inconsistency. Their method, also known as AFSfMlearner, comprises four modules. The Structure module (DepthNet) uses a ResNet-18 encoder-decoder architecture to estimate dense depth maps. The Motion module (PoseNet) is a lightweight network that predicts six Degree of freedom relative camera poses between frames. The Appearance module predicts appearance flow and calibrates brightness, using a structure similar to DepthNet. The Correspondence module performs automatic registration to enhance performance, with architecture similar to the appearance module. The appearance flow allows relaxation of the brightness constancy assumption, addressing severe brightness fluctuations common in endoscopic scenarios.
OneSLAM [^4] https://github.com/arcadelab/OneSLAM [25] presents a generalized SLAM approach for monocular endoscopic imaging based on tracking any point. It leverages a tracking any point foundation model to obtain sparse correspondences across multiple frames. These are used to perform local bundle adjustment to jointly optimize camera poses and 3D structure. OneSLAM demonstrates strong cross-domain performance on sinus endoscopy, colonoscopy, arthroscopy and laparoscopy data without retraining.
All competitor methods were utilized as developed by their respective authors, without modifications. This varied baseline allows us to comprehensively assess the generalization capabilities and performance of our proposed MVSLAM approach across different endoscopic scenarios and datasets.
@@ 5.1 Performance Metrics and Statistical Analysis
To evaluate the performance of the SLAM approach, we employed a set of well-established metrics for depth estimation and pose estimation. For depth estimation, we selected the Absolute Relative Difference, Root Mean Square Error (RMSE), RMSE log, Squared Relative Error, and accuracy thresholds of $1.25$, $1.25^{2}$, and $1.25^{3}$, as described in [42] and presented in Table 2.
@@
Table 3: Comparative Overview of Trajectory Evaluation Metrics. $E_{ij}$represents the pose error between estimated and ground truth poses. The function trans($E_{ij}$) extracts the translational component, and $angle(\cdot)$ calculates the angular difference from rotational data.
To measure the pose estimation performance, we utilized the Relative Translation Error (RTE), Relative Rotation Error (RRE), and Absolute Translation Error (ATE), as shown in Table 3. These metrics are widely used in the literature for evaluating the accuracy of pose estimation in SLAM systems [1].
In addition to the performance metrics, we conducted a statistical analysis to determine the significance of the performance differences observed between the MVSLAM approach and the benchmark models across various scenarios. We verified that the data for each model meets the assumptions required for using an independent two-sample T-test, including normality and equal variances. We then performed pairwise comparisons between the MVSLAM approach and each benchmark model separately, using a significance level of $p=0.05$.
@@ 6 Results
In this section, we present the results of our comprehensive evaluation aimed at validating the efficacy and robustness of our proposed MVSLAM approach. Initially, in section 6.1, we provide an analysis of the monocular pose estimation module. This is followed by section 6.2, where we explore the application and validation of the depth estimation module. It’s worth noting that our preprocessing methods for certain datasets differ from some competitors, which may lead to variations in results. For instance, we used a version of the Hamlyn dataset that had been preprocessed using libellas, as this preprocessed version has been utilized in other notable works such as [22]. This choice ensures consistency and comparability with existing literature in the field. For the SCARED Dataset, we derived depth information through stereo rectification and disparity estimation, as opposed to extracting it directly from the point cloud as some other studies have done. We opted for this approach because it allows us to obtain a dense depth map, in contrast to the sparse depth information typically yielded by point cloud extraction. It should be noted that the 3D Mapping Module (3DM) cannot be directly evaluated due to two key factors. First, the lack of ground truth 3D models for the tested environments hinders a comprehensive validation of the reconstruction quality. Second, even if ground truth 3D models were available, the scale ambiguity inherent in the monocular pose and depth estimation modules would render the reconstructed model incomparable due to differing scales. However, this inability to directly evaluate the 3DM does not represent a limitation of our approach. The 3DM employs state-of-the-art 3D reconstruction algorithms that have been rigorously tested and validated in prior work [39, 43]. Furthermore, the 3D mapping and localization capabilities critically depend on the performance of the pose estimation and depth estimation modules, which are thoroughly evaluated in the following sections. The successful functioning of the overall SLAM system thus implicitly validates the efficacy of the 3DM.
@@ 6.1 Monocular Pose Estimation Module (MPEM)
Figure 4 presents the performances of the monocular pose estimation module measured in terms of ATE, RTE, and RRE on the two datasets (SCARED and EndoSLAM) when adopting each of the five methods, i.e., EndoSfmLearner, EndoDepth, AFSfMlearner, OneSLAM and CycleVO. When interpreting these results, it’s important to consider the training backgrounds of the methods. EndoSLAM was trained on the EndoSLAM dataset, while AFSfMLearner was trained on the SCARED dataset. This prior exposure to the test data or related datasets may introduce a potential bias in their performance, which should be taken into account when interpreting their results. In contrast, CycleVO has not been exposed to either dataset during its training.
Figure 4: Comparison of the performance of EndoDepth, EndoSfmLearner, AFSfMlearner, OneSLAM and CycleVO algorithms on the SCARED and ENDOSLAM datasets. Metrics evaluated include Absolute Trajectory Error (ATE), Relative Trajectory Error (RTE), and Relative Rotation Error (RRE). Each box plot shows the distribution of the errors for the respective algorithms, highlighting the median, interquartile range, and outliers.
In the case of trajectory errors, being either an absolute or a relative error, both rows show that CycleVO and EndoSfmLearner have lower median scores and interquartile ranges than EndoDepth, AFSfMlearner and OneSLAM. The figure compares the three methods across the two datasets and three metrics. The rows represent the SCARED and ENDOSLAM datasets, while the columns show the ATE, RTE, and RRE metrics respectively. Regarding the statistical analysis, Table 4 shows the pairwise comparisons between CycleVO and the other methods. Notably, CycleVO achieves competitive performance without any prior exposure to either the EndoSLAM or SCARED datasets during its training, highlighting its strong generalization capabilities. It is also interesting to note that our CycleVO approach not only performs as the others but it also has the lowest inference time (Table 5).
Metric Comparison T-Stat P-Value
@@
BodySLAM vs EndoSfMLearner 0.63 0.5335
BodySLAM vs OneSLAM -11.79 1.81e-10
BodySLAM vs AFSfMlearner -7.01 1.20e-06
@@
BodySLAM vs OneSLAM -9.00 3.70e-08
BodySLAM vs AFSfMlearner 4.17 0.0002
ENDOSLAM ATE BodySLAM vs EndoDepth -1.02 0.3122
BodySLAM vs EndoSfMLearner -0.11 0.9087
BodySLAM vs OneSLAM -0.34 0.7343
BodySLAM vs AFSfMlearner -2.68 0.0090
@@
BodySLAM vs EndoSfMLearner -0.01 0.9942
BodySLAM vs OneSLAM -9.79 7.24e-14
BodySLAM vs AFSfMlearner -7.60 1.73e-10
@@
BodySLAM vs OneSLAM -15.39 5.06e-20
BodySLAM vs AFSfMlearner -0.63 0.5299
Table 4: Comparative analysis of BodySLAM against EndoDepth, EndoSfMLearner, OneSLAM, and AFSfMlearner across different metrics (ATE, RTE, RRE) for SCARED and ENDOSLAM datasets. T-Statistic and P-Value are reported for each comparison. Bold P-Values indicate statistically significant differences (p < 0.05).
Model Dataset Average Time [s]
EndoDepth EndoSLAM 1140.56
SCARED 934.18
@@
AFSfMlearner EndoSLAM 14.63
SCARED 59.95
OneSLAM EndoSLAM 169.77
SCARED 336.65
CycleVO EndoSLAM 3.32
@@
Absolute Relative Difference BodySLAM vs EndoDepth -2.18 0.0366
BodySLAM vs EndoSfMLearner -4.24 0.0002
BodySLAM vs AFSfMLearner -12.50 3.55e-13
RMSE BodySLAM vs EndoDepth -1.99 0.0550
@@
BodySLAM vs AFSfMLearner -6.26 2.73e-06
RMSE Log BodySLAM vs EndoDepth -2.13 0.0400
BodySLAM vs EndoSfMLearner -4.42 9.94e-05
BodySLAM vs AFSfMLearner -17.08 1.29e-15
Squared Relative Error BodySLAM vs EndoDepth -2.15 0.0422
BodySLAM vs EndoSfMLearner -2.71 0.0105
BodySLAM vs AFSfMLearner -4.80 0.0001
accuracy (1.25) BodySLAM vs EndoDepth 5.14 4.42e-05
BodySLAM vs EndoSfMLearner 2.55 0.0151
BodySLAM vs AFSfMLearner 14.17 2.55e-13
accuracy (1.25)² BodySLAM vs EndoDepth 4.12 0.0006
BodySLAM vs EndoSfMLearner 1.54 0.1396
BodySLAM vs AFSfMLearner 10.38 4.62e-09
accuracy (1.25)³ BodySLAM vs EndoDepth 3.56 0.0022
BodySLAM vs EndoSfMLearner 1.37 0.1886
BodySLAM vs AFSfMLearner 14.10 3.62e-11
Table 6: Statistical Performance Analysis of the Monocular Depth Estimation Module on the Hamlyn Dataset. The table details outcomes for Absolute Relative Difference, Root Mean Square Error (RMSE), RMSE Log, Squared Relative Error, and accuracy thresholds (accuracy < 1.25, accuracy < 1.25², accuracy < 1.25³). Results include pairwise comparisons of BodySLAM versus EndoDepth, EndoSfMLearner, and AFSfMLearner, assessed through t-tests with a significance level of p < 0.05.
@@
Absolute Relative Difference BodySLAM vs EndoDepth -9.69 1.81e-09
BodySLAM vs EndoSfMLearner -4.14 1.99e-04
BodySLAM vs AFSfMlearner -4.36 3.39e-04
RMSE BodySLAM vs EndoDepth -1.01 0.3212
@@
BodySLAM vs AFSfMlearner -0.78 0.4392
RMSE Log BodySLAM vs EndoDepth -14.99 2.72e-13
BodySLAM vs EndoSfMLearner -8.35 1.05e-09
BodySLAM vs AFSfMlearner -13.13 4.65e-11
Squared Relative Error BodySLAM vs EndoDepth -0.85 0.4069
BodySLAM vs EndoSfMLearner 0.31 0.7601
BodySLAM vs AFSfMlearner -0.35 0.7305
accuracy (1.25) BodySLAM vs EndoDepth 13.10 3.61e-11
BodySLAM vs EndoSfMLearner 5.82 6.53e-06
BodySLAM vs AFSfMlearner 9.36 1.69e-08
accuracy (1.25)² BodySLAM vs EndoDepth 10.13 7.29e-09
BodySLAM vs EndoSfMLearner 2.56 0.0195
BodySLAM vs AFSfMlearner 6.26 6.59e-06
accuracy (1.25)³ BodySLAM vs EndoDepth 8.53 9.75e-08
BodySLAM vs EndoSfMLearner 1.26 0.2228
BodySLAM vs AFSfMlearner 5.47 3.40e-05
Table 7: Statistical Performance Analysis of the Monocular Depth Estimation Module on the SCARED Dataset. The table details outcomes for Absolute Relative Difference, Root Mean Square Error (RMSE), RMSE Log, Squared Relative Error, and accuracy thresholds (accuracy < 1.25, accuracy < 1.25², accuracy < 1.25³). Results include pairwise comparisons of BodySLAM versus EndoDepth, EndoSfMLearner, and AFSfMlearner, assessed through t-tests with a significance level of p < 0.05.
Turning our attention to RRE, i.e., a score measuring the angular deviation in the rotation component, we notice that CycleVO outperforms both EndoSfmLearner, EndoDepth, AFSfMlearner and OneSLAM on both datasets (both rows and last column of Figure 4). Furthermore, these performance differences are also statistically significant (last rows for both datasets of Figure 4). To deepen the results we also perform an ablation study that aims to investigate the effectiveness of the new terms that we introduced for the cycle loss (section 3.1). To this end, we ran four experiments, two per dataset, setting $\lambda_{2}\neq 0$, $\lambda_{2}=0$ (Figure 5).